# Synthetic Prior for Few-Shot Drivable Head Avatar Inversion

### v2

@@ 5 Discussion
While significantly outperforming monocular methods, SynShot has certain limitations that we identify. A key challenge is bridging the domain gap between synthetic and real data. There is considerable room for improvement in the generation of synthetic data. For instance, currently, all our synthetic subjects share the same teeth geometry and texture. As a consequence, teeth in our inverted head avatars often closely follow the prior and do not adapt easily. Furthermore, our synthetic data lacks diverse expression-dependent wrinkles, affecting its overall visual quality. Additionally, our dataset was ray-traced with a single environment map, limiting generalization to varied lighting conditions. A key challenge for future work is to increase the diversity of the synthetic dataset, thereby improving the visual quality of the reconstructed head avatars.
@@ 6 Conclusion
We have proposed SynShot, a method for reconstructing a personalized 3D Gaussian head avatar from just a few images. SynShot builds a generative head avatar purely from synthetic data and then utilizes this model as a prior in an inversion pipeline. This inversion pipeline follows a pivotal tuning strategy that successfully bridges the domain gap between the prior and the real input images. We demonstrate that our personalized head avatars generalize better to unseen expressions and viewpoints than SOTA personalized head avatars. This demonstrates the utility of synthetic data for creating photo-realistic 3D head avatars. Our method has numerous potential downstream applications, including holoportation, self and cross-reenactment, and virtual presence. Furthermore, we can distill the generated Gaussian point cloud into a lightweight, network-free representation using GEM [88], further extending SynShot’s capabilities and removing the need for a high-end GPU.
@@ Appendix A Appendix
All our inversion results used only three input images (Figure 22) unless stated otherwise. Figure 17 compares monocular baseline methods trained on the entire dataset with our inversion approach. Furthermore, we provide additional examples of cross-reenactment comparisons, demonstrating the advantages of our method compared to baselines trained on only 13 frames. Next, we present results with progressively varying numbers of training frames, illustrating how this influences the quality of reconstruction. Figures 18 and 19 highlight the importance of our synthetic prior.

### v3

@@ 5 Discussion
While significantly outperforming monocular methods, SynShot has certain limitations that we identify. A key challenge is bridging the domain gap between synthetic and real data. There is considerable room for improvement in the generation of synthetic data. For example, all our synthetic subjects share the same teeth geometry and texture. As a consequence, teeth in our inverted head avatars often closely follow the prior and do not adapt easily. Furthermore, our synthetic data lacks diverse expression-dependent wrinkles, affecting its overall visual quality. Additionally, our dataset was ray-traced with a single environment map, limiting generalization to varied lighting conditions.
@@ 6 Conclusion
We have proposed SynShot, a method for reconstructing a personalized 3D Gaussian head avatar from just a few images. SynShot builds a generative head avatar purely from synthetic data and then utilizes this model as a prior in an inversion pipeline. This inversion pipeline follows a pivotal tuning strategy that successfully bridges the domain gap between the prior and the real input images. We demonstrate that our personalized head avatars generalize better to unseen expressions and viewpoints than SOTA head avatars.
@@ Appendix A Appendix
All our inversion results used only three input images (Figure 21) unless stated otherwise. Figure 16 compares monocular baseline methods trained on the entire dataset with our inversion approach. Furthermore, we provide additional examples of cross-reenactment comparisons, demonstrating the advantages of our method compared to baselines trained on only 13 frames. Next, we present results with progressively varying numbers of training frames, illustrating how this influences the quality of reconstruction. Figures 17 and 18 highlight the importance of our synthetic prior.